# Two-Stage Learning of Stabilizing Neural Controllers via Zubov Sampling and Iterative Domain Expansion

### v1

@@ Abstract
Learning-based neural network (NN) control policies have shown impressive empirical performance. However, obtaining stability guarantees and estimations of the region of attraction of these learned neural controllers is challenging due to the lack of stable and scalable training and verification algorithms. Although previous works in this area have achieved great success, much conservatism remains in their framework. In this work, we propose a novel two-stage training framework to jointly synthesize the controller and Lyapunov function for continuous-time systems. By leveraging a Zubov-inspired region of attraction characterization to directly estimate stability boundaries, we propose a novel training data sampling strategy and a domain updating mechanism that significantly reduces the conservatism in training. Moreover, unlike existing works on continuous-time systems that rely on an SMT solver to formally verify the Lyapunov condition, we extend state-of-the-art neural network verifier $\alpha,\!\beta$-CROWN with the capability of performing automatic bound propagation through the Jacobian of dynamical systems and a novel verification scheme that avoids expensive bisection. To demonstrate the effectiveness of our approach, we conduct numerical experiments by synthesizing and verifying controllers on several challenging nonlinear systems across multiple dimensions. We show that our training can yield region of attractions with volume $5-1.5\cdot 10^{5}$ times larger compared to the baselines, and our verification on continuous systems can be up to $40-10000$ times faster compared to the traditional SMT solver dReal. Our code is available at https://github.com/Verified-Intelligence/Two-Stage_Neural_Controller_Training.
@@ 1 Introduction
Learning-based neural network control policies have demonstrated great potential in complex systems due to their high expressiveness (zhang2016learning, ; wang2020reinforcement, ; kaufmann2023champion, ; zhao2024applications, ). However, the applications of these neural controllers in safety-critical real-world physical systems are of concern since they typically lack stability guarantees. A promising method for quantifying safety is estimating the region of attraction (ROA), which refers to the set of initial states from which a system is guaranteed to converge to a desired equilibrium point. Computation of the ROA is often formulated as the search for an appropriate Lyapunov function satisfying algebraic conditions over its sub-level sets (lyapunov1992general, ; valmorbida2017region, ; ratschan2010providing, ). Due to the highly complex and nonlinear nature of neural networks, searching for a Lyapunov function is not easy. Many existing works that are able to make stability guarantees about a system typically have some restrictions on the problem structure, for example, requiring linear or polynomial parameterizations of the system, where sum-of-squares technique (tedrake2010lqr, ; yang2023approximate, ; papachristodoulou2002construction, ; dai2023convex, ) can be applied. To address this limitation and enable certification for general nonlinear systems, many recent works explore data-driven methods to synthesize a neural certificate, such as Lyapunov functions (chang2019neural, ; dai2021lyapunov, ; zhou2022neural, ; wu2023neural, ; yang2024lyapunov, ; shi2024certified, ; liu2025physics, ), Barrier Functions (ames2016control, ; robey2020learning, ; dai2023convex, ; dawson2022safe, ; liu2023safe, ; dawson2022learning, ), and Contraction Metrics (tsukamoto2020neural, ; tsukamoto2020stochastic, ; tsukamoto2021theoretical, ; dawson2023safe, ; sun2021learning, ), to certify the system’s stability.
@@
In this work, we aim to synthesize neural controllers of continuous-time systems together with a Lyapunov certificate. Unlike typical machine learning tasks, this problem requires that the Lyapunov and boundary conditions hold over the entire continuous domain of interest, rather than statistically achieving high accuracy on a discrete test set. However, since at each iteration only a small number of points can be selected, the problem forces us to adopt specialized strategies both to pick the training domain and to sample training data from the domain. We argue that these two are the key factors for the conservativeness issue presented in many previous works. For the training domain selection, many previous works simply select a fixed domain of interest and enforce the Lyapunov condition over it. While some previous works proposed a new formulation to overcome the conservativeness by introducing an extra term to verify forward invariance and allow the ROA to touch the training domain non-trivially (yang2024lyapunov, ; shi2024certified, ), we argue that this does not fully solve the conservativeness issue and leaves the training domain selection a hard hyperparameter to tune.
@@
In addition to the training domain, training data selection also plays an important role. Besides simple random sampling that typically does not suffice for certification, previous works dealing with similar problems (chang2019neural, ; dai2021lyapunov, ; zhou2022neural, ; yang2024lyapunov, ) typically achieve co-learning with a counterexample-guided (CEGIS) approach, which generates counterexamples of the Lyapunov condition either using a verifier (chang2019neural, ; dai2021lyapunov, ) such as an SMT or MILP solver or a gradient-based approach (wu2023neural, ; yang2024lyapunov, ) to be the training data. While these approaches have led to great success, such approaches can typically only refine upon already worked controller/Lyapunov function pair and thus rely on carefully tuned LQR/RL based initializations (chang2019neural, ; wu2023neural, ; yang2024lyapunov, ). This limitation makes these works both conservative in their estimation of the ROA and also very difficult to apply to new systems without much domain-specific knowledge.
@@
In our work, we propose a two-stage training framework with carefully designed curriculum domain expansion and training sample selection that can lead to much more non-conservative ROA estimation inspired by Zubov theorem (zubov1961methods, ; liu2025physics, ; kang2023data, ; meng2024zubov, ),which provides a characterization of the true ROA with some PDE. During the first stage, we propose a Zubov-inspired training data sampling strategy by combining points from inside and outside of the current ROA estimation to better guide the learning. Moreover, we eliminate the conservativeness introduced by the training domain by dynamically expanding it using trajectory information. This stage can provide us an almost working controller and Lyapunov function that we can refine upon. The second stage then involves counterexample-guided learning (CEGIS) to eliminate all the counterexamples inside the ROA estimation and thus obtain a verifiable Lyapunov certificate. As a by-product, our training also fully utilizes the physics-informed Zubov loss and can eliminate the conservatism introduced by the inner approximation nature of the normal Lyapunov function. As we shall demonstrate in our numerical experiments, our approach can robustly produce a controller with a significantly larger ROA compared to previous state-of-the-art approaches in all the benchmarks with various control input limits.
@@
As we deal with continuous-time systems, verification over Jacobian vector products also presents its challenge. Previous works addressing the Lyapunov stability verification for continuous-time systems typically rely on SMT or SDP based solvers (chang2019neural, ; abate2020formal, ; liu2025physics, ; chen2021learning, ; yin2021stability, ; fazlyab2020safety, ; chen2021learning, ) such as dReal (gao2013dreal, ). Inspired by the success in (yang2024lyapunov, ; shi2024certified, ), we also adapt the state-of-the-art neural network verification tool $\alpha,\!\beta$-CROWN (zhang2018efficient, ; salman2019convex, ; xu2020automatic, ; xu2021fast, ; wang2021beta, ; zhang2022general, ) for formally verifying Lyapunov stability. The key challenge is to handle the Jacobian-related operators in dynamical systems, which have not been discussed or implemented in existing neural network verifiers. To ensure verification tightness, we design new linear relaxations for these Jacobian operators in continuous time system learning such as Tanh and Sigmoid. Furthermore, we introduced a novel verification scheme to dynamically adjust the specifications without incurring much computational cost to avoid expensive bisections as in (yang2024lyapunov, ; liu2025physics, ). To summarize, our main contribution includes:
@@
  • We customize and strengthen advanced neural network verification tool $\alpha,\!\beta$-CROWN with the ability of handling many commonly used Jacobian-related operators in continuous-time systems. Moreover, we propose a novel verification algorithm that can effectively avoid expensive bisections without sacrificing on soundness.
  • Through numerical experiments, we demonstrate that our training can yield region of attraction with volume $5-1.5\cdot 10^{5}$ times larger than the baselines. Moreover, across the systems under evaluation which we can formally verify, our verification achieves $40-10000$ times speedup compared to the previous SMT solver dReal.
@@ 2 Background and Problem Statement
where $x$ is the system state, $g$ denotes the open-loop dynamics, $u$ denotes the controller, and $f$ denotes the closed-loop nonlinear dynamics. We denote the solution of the system starting with initial condition $x=x_{0}$ by $\phi(t,x_{0})$, and denote the equilibrium state and the control input as $x^{*}/u^{*}$ respectively. We denote the sublevelset $\{x:f(x)\leq c\}$ of some function $f$ as $f^{\leq c}$ and correspondingly $f^{=c}$ for the levelset. Our objective is to jointly search a neural network control policy $u_{\theta}$ together with a Lyapunov function $V_{\theta}$ that can certify the stability of the closed-loop dynamics (1) according to the Lyapunov theorem. The goal is to learn a controller that maximize the region of attraction (ROA), which has the formal definition below.
@@
Region of attraction for the system (1) is the set $\mathcal{R}$ such that $\lim_{t\to\infty}\phi(t,x_{0})=x^{*}$ for all $x_{0}\in\mathcal{R}$.
@@
$V(x^{\!*})$ $=0,$ (2)
$V(x)$ $>0\;\;(\forall\,x\in\mathcal{S}\setminus\{x^{\!*}\}),$
$\nabla V(x)^{\!\top}f(x)$ $<0\;\;(\forall\,x\in\mathcal{S}),$
@@
To maximize the possible ROA, we adopt the Zubov theorem, which states that a function satisfies a stricter set of constraints than the Lyapunov condition can express the real underlying ROA $\mathcal{R}$.
@@
Consider the nonlinear dynamical system in (1) and we WLOG assume that $x^{*}=0$. Let $\mathcal{A}\subset\mathbb{R}^{n}$ and assume that there exists a continuously differentiable function $W:\mathcal{A}\to\mathbb{R}$ and a positive definite function $\Psi:\mathbb{R}^{n}\to\mathbb{R}$ satisfying the following conditions,
@@
Then $x^{*}=0$ is asymptotically stable with the region of attraction exactly equals to $\mathcal{A}$.
@@
[…]
$W(x)=\begin{cases}\beta(V(x)),\quad&\text{if }V(x)<\infty\\ 1,\quad&\text{otherwise}\end{cases}.$ (7)
[…]
$\nabla W(x)^{\top}f(x)=-a(1+W(x))(1-W(x))\|x\|^{p}.$ (8)
We refer readers to liu2025physics for more theoretical details. This theorem allows us to obtain the ground truth value of $W$ during training without relying solely on the PDE residual loss, which makes the training process more robust and can lead to improved learning dynamics for the Lyapunov function.
@@ 3.1 Learning of Neural Controller and Lyapunov Function
$u_{\theta}(x)$ $=c\cdot\tanh(\text{NN}_{c}(x)-\text{NN}_{c}(x^{*})+\tanh^{-1}(\frac{u^{*}}{c}))$ (9)
$V_{\theta}(x)$ $=\text{sigmoid}(\text{NN}_{v}(x))$ (10)
This parametrization ensures that we have $0\leq V_{\theta}\leq 1$ and the crucial property $u_{\theta}(0)=u^{*}$, which ensures that the origin is the equilibrium also for the closed-loop system. Moreover, this parametrization constrains the input to be in the range $[-c,c]$. Sometimes, if a non-symmetric limit $[0,c]$ is more natural, we add another layer of ReLU to $u_{\theta}$, and this operation can be applied to each coordinate to support a mixed control input limit.
@@ 3.1.1 ROA Estimation Stage
The first stage of colearning which we call ROA estimation stage, is aiming to set up a mostly working controller and an estimation of Lyapunov function that can be further refined. As we have discussed, in this stage the most important aspect is the selection of training data and training domain.
Zubov Guided Sampling:
@@
Intuitively, a good training data selection scheme should mix the data points inside and outside of the running region of attraction. Selecting only “outside” points risks collapsing the sublevel set back toward the origin, while sampling exclusively “inside” offers no signal to enlarge it. In low dimensions, a well-chosen box plus random sampling can approximate this mix, but as dimensionality grows, the ROA occupies a vanishing fraction of the domain, and naive draws will almost always only select the points outside of the ROA. For example, in the cartpole system, the ROA we found at the end of training only occupies less than $1\%$ of the full training domain, and this will only be more pronounced in even higher-dimensional systems, and therefore naive sampling is infeasible.
In the traditional Lyapunov function framework, it is difficult to obtain this running estimation of the ROA. Zubov theorem, on the contrary, characterizes the exact region of attraction by the sublevel set $V^{\leq 1}$. Intuitively, In our training pipeline, since our parametrization already deems that $V$ cannot exactly equal $1$, we propose picking the training data by selecting a batch of points lying in $V_{\theta}^{\leq c}$ and another batch with $V_{\theta}(x)$ as close to $1$ as possible with $c\approx 1$. To achieve this, we select data by performing projected gradient descent (PGD) madry2017towards with the following two objectives
@@
The first interior loss will push points to the sublevel set $V_{\theta}^{\leq c}$ and the outside loss will push points further away from the region of attraction as shown in figure 1.
@@
Dynamic Training Domain Expansion: While previous works attempt to resolve the conservativeness introduced by training domain by allowing the ROA to touch non-trivially with it, we argue that the best way is instead to dynamically expand the domain during training. We thus propose to utilize the trajectory information to dynamically update the training box. We start from a relatively small training region and periodically perform a trajectory simulation starting from the points inside the current ROA estimation $V_{\theta}^{\leq c}$ with the same $c$ in the data selection. We then expand the training region to include the farthest states that the converged trajectories can reach. To also benefit simple systems where all trajectories converge to zero in a monotone fashion, we further enlarge the box uniformly for the final training domain. If there is no convergent trajectory for many iterations, we manually enlarge the box for further exploration. This approach simultaneously resolves conservativeness in ROA posed by fixed domain and also resolves the challenging forward invariance issue. A visualization of our approach can be found in figure 1, where the original blue training domain is expanded further to incorporate the states along the convergent trajectories. The domain updating algorithm can be found in algorithm 2 in appendix.
[…]
@@
With this approach, only simulation to time $T$ is required. In practice, we found that this term can benefit training even if a very small $T$ like $T=0.01s$ is adopted. We also utilized a similar strategy as in (wang2024actor, ) to decouple the learning of the Lyapunov function and controller using the actor-critic framework and impose the boundary condition $V_{\theta}(x)\approx 1$ for $x\in\partial(2\Omega)$ to guide the early phase learning. The full loss function can be found in appendix section A.3.
3.1.2 CEGIS based Refinement

### v2

# Two‑Stage Learning of Stabilizing Neural Controllers via Zubov Sampling and Iterative Domain Expansion

@@ Abstract
Learning-based neural network (NN) control policies have shown impressive empirical performance. However, obtaining stability guarantees and estimates of the region of attraction of these learned neural controllers is challenging due to the lack of stable and scalable training and verification algorithms. Although previous works in this area have achieved great success, much conservatism remains in their frameworks. In this work, we propose a novel two-stage training framework to jointly synthesize a controller and a Lyapunov function for continuous-time systems. By leveraging a Zubov‑inspired region of attraction characterization to directly estimate stability boundaries, we propose a novel training-data sampling strategy and a domain-updating mechanism that significantly reduces the conservatism in training. Moreover, unlike existing works on continuous-time systems that rely on an SMT solver to formally verify the Lyapunov condition, we extend state-of-the-art neural network verifier $\alpha,\!\beta$-CROWN with the capability of performing automatic bound propagation through the Jacobian of dynamical systems and a novel verification scheme that avoids expensive bisection. To demonstrate the effectiveness of our approach, we conduct numerical experiments by synthesizing and verifying controllers on several challenging nonlinear systems across multiple dimensions. We show that our training can yield region of attractions with volume $5-1.5\cdot 10^{5}$ times larger compared to the baselines, and our verification on continuous systems can be up to 4010,00040-10{,}000 times faster compared to the traditional SMT solver dReal. Our code is available at https://github.com/Verified-Intelligence/Two-Stage_Neural_Controller_Training.
@@ 1 Introduction
Learning-based neural network control policies have demonstrated great potential in complex systems due to their high expressiveness (kaufmann2023champion, ; zhao2024applications, ). However, the applications of these neural controllers in safety-critical real-world physical systems raise concerns since they typically lack stability guarantees. A promising method for quantifying safety is estimating the region of attraction (ROA), which refers to the set of initial states from which a system is guaranteed to converge to a desired equilibrium point. The computation of the ROA is often formulated as the search for an appropriate Lyapunov function satisfying some algebraic conditions over its sublevel sets (lyapunov1992general, ; valmorbida2017region, ; ratschan2010providing, ). Due to the highly complex and nonlinear nature of neural networks, searching for a Lyapunov function is not easy. Many existing works that are able to make stability guarantees about a system have some restrictions on the problem structure, for example, requiring linear or polynomial parameterizations of the system, where sum-of-squares technique (tedrake2010lqr, ; yang2023approximate, ; papachristodoulou2002construction, ; dai2023convex, ) can be applied. To enable certification for general nonlinear systems, many recent works explore data-driven methods to synthesize a neural certificate (dawson2023safe, ), such as Lyapunov functions (chang2019neural, ; dai2021lyapunov, ; zhou2022neural, ; wu2023neural, ; yang2024lyapunov, ; shi2024certified, ; liu2025physics, ; serry2025safe, ; meng2025learning, ; liu2024tool, ), barrier Functions (ames2016control, ; robey2020learning, ; dai2023convex, ; dawson2022safe, ; liu2023safe, ; dawson2022learning, ; jagtap2020control, ; jagtap2020formal, ), and contraction metrics (tsukamoto2020neural, ; tsukamoto2020stochastic, ; tsukamoto2021theoretical, ; sun2021learning, ; li2025neural, ), to certify the system’s stability.
@@
In this work, we aim to synthesize neural controllers of continuous-time systems together with a Lyapunov certificate. Unlike typical machine learning tasks, this problem requires that the Lyapunov and boundary conditions hold over the entire continuous domain of interest, rather than statistically achieving high accuracy on a discrete test set. However, since at each iteration only a small number of points can be selected, the problem forces us to adopt specialized strategies both to pick the training domain and to sample training data from the domain. We argue that these two are the key factors for the conservativeness issue observed in many previous works. For selecting the training domain, many previous works simply select a fixed domain of interest and enforce the Lyapunov condition over it. While some previous works have proposed a new formulation to overcome the conservativeness by introducing an extra term to verify forward invariance and allow the ROA to touch the training domain nontrivially (yang2024lyapunov, ; shi2024certified, ), we argue that this does not fully solve the conservativeness issue and leaves the training domain selection a hard hyperparameter to tune.
@@
In addition to the training domain, training data selection also plays an important role. Besides simple random sampling that typically does not suffice for certification, previous works dealing with similar problems (chang2019neural, ; dai2021lyapunov, ; zhou2022neural, ; yang2024lyapunov, ) typically achieve co-learning with a counterexample-guided (CEGIS) approach, which generates counterexamples of the Lyapunov condition as the training data, either using a verifier (chang2019neural, ; dai2021lyapunov, ) such as an SMT or MILP solver or a gradient-based approach (wu2023neural, ; yang2024lyapunov, ). While these approaches have led to great success, such approaches can typically only refine upon already worked controller/Lyapunov function pair and thus rely on carefully tuned LQR-/RL-based initializations (chang2019neural, ; wu2023neural, ; yang2024lyapunov, ). This limitation makes these works both conservative in their estimation of the ROA and also very difficult to apply to new systems without much domain-specific knowledge.
@@
In our work, we propose a two-stage training framework with carefully designed curriculum domain expansion and training sample selection that can yield significantly less conservative ROA estimation inspired by Zubov theorem (zubov1961methods, ; liu2025physics, ; kang2023data, ; meng2024zubov, ), which provides a characterization of the true ROA with a PDE. During the first stage, we propose a Zubov-inspired training data sampling strategy by combining points from both inside and outside of the current ROA estimation to better guide the learning. Moreover, we eliminate the conservativeness introduced by the training domain by dynamically expanding it using trajectory information. This stage can provide us an almost-working controller and Lyapunov function that we can refine upon. The second stage then involves counterexample-guided learning (CEGIS) to eliminate all the counterexamples inside the ROA estimation and thus obtain a verifiable Lyapunov certificate. Importantly, our training also fully utilizes the physics-informed Zubov loss and can eliminate the conservatism introduced by the inner approximation nature of the normal Lyapunov function. As we shall demonstrate in our numerical experiments, our approach can robustly produce controllers with significantly larger ROAs compared to previous state-of-the-art approaches in all the benchmarks with various control input limits.
@@
As we deal with continuous-time systems, verification over Jacobian vector products also presents a challenge. Previous works addressing the Lyapunov stability verification for continuous-time systems typically rely on SMT-, MIP-, or SDP-based solvers (chang2019neural, ; abate2020formal, ; liu2025physics, ; chen2021learning, ; yin2021stability, ; fazlyab2020safety, ; chen2021learning, ) such as dReal (gao2013dreal, ) or Z3 (de2008z3, ). Inspired by the success in (yang2024lyapunov, ; shi2024certified, ), we also utilize the state-of-the-art neural network verification tool $\alpha,\!\beta$-CROWN (zhang2018efficient, ; salman2019convex, ; xu2020automatic, ; xu2021fast, ; wang2021beta, ; zhang2022general, ) for formally verifying Lyapunov stability. The key challenge is to handle the Jacobian-related operators in dynamical systems, which have not been discussed or implemented in existing neural network verifiers. To ensure verification tightness, we design new linear relaxations for these Jacobian operators in continuous time system learning such as Tanh and Sigmoid. Furthermore, we introduce a novel verification scheme to dynamically adjust the specifications when noticing counterexamples without incurring much computational cost to avoid expensive bisections as in (yang2024lyapunov, ; liu2025physics, ).
To summarize, our main contributions include:
@@
  • We customize and strengthen the advanced neural network verification tool $\alpha,\!\beta$-CROWN with the ability of handling many commonly used Jacobian-related operators in continuous-time systems. Moreover, we propose a novel verification algorithm that can effectively avoid expensive bisections without sacrificing on soundness.
  • We demonstrate through numerical experiments that our training can yield region of attraction with volume $5-1.5\cdot 10^{5}$ times larger than the baselines. Moreover, across the systems under evaluation which we can formally verify, our verification achieves 4010,00040-10{,}000 times speedup compared to the previous SMT solver dReal.
@@ 2 Background and Problem Statement
where $x$ is the system state, $g$ denotes the open-loop dynamics, $u$ denotes the controller, and $f$ denotes the closed-loop nonlinear dynamics. We denote the solution of the system starting with initial condition $x=x_{0}$ by $\phi(t,x_{0})$, and denote the equilibrium state and the control input as $x^{\ast}/u^{\ast}$ respectively. We denote the sublevelset $\{x:f(x)\leq c\}$ of some function $f$ as $f^{\leq c}$ and correspondingly $f^{=c}$ for the levelset. Our objective is to jointly search for a neural network control policy $u_{\theta}$ together with a Lyapunov function $V_{\theta}$ that can certify the stability of the closed-loop dynamics (1) according to the Lyapunov theorem. The goal is to learn a controller that maximizes the region of attraction (ROA), which has the formal definition below.
@@
Region of attraction for the system (1) is the set $\mathcal{R}$ such that $\lim_{t\to\infty}\phi(t,x_{0})=x^{\ast}$ for all $x_{0}\in\mathcal{R}$.
@@
$V(x^{\ast})$ $=0$ (2)
$V(x)$ $>0\;\;(\forall\,x\in\mathcal{S}\setminus\{x^{\ast}\})$
$\nabla V(x)^{\!\top}f(x)$ $<0\;\;(\forall\,x\in\mathcal{S})$
@@
To maximize the possible ROA, we adopt the Zubov theorem, which states that a function that satisfies a stricter set of constraints than the Lyapunov condition can express the real underlying ROA $\mathcal{R}$.
@@
Consider the nonlinear dynamical system in (1) and we assume that $x^{\ast}=0$ without loss of generality. Let $\mathcal{A}\subset\mathbb{R}^{n}$ and assume that there exists a continuously differentiable function $W:\mathcal{A}\to\mathbb{R}$ and a positive definite function $\Psi:\mathbb{R}^{n}\to\mathbb{R}$ satisfying the following conditions,
@@
Then $x^{\ast}=0$ is asymptotically stable with the region of attraction is exactly $\mathcal{A}$.
@@
We adopt the convention that divergence of integration yields the value $\infty$. Then for any $a,p>0$, $W(x)=\tanh(a\int_{0}^{\infty}\|\phi(t,x)\|^{p}\,dt)$ solves the Zubov equation equation $\nabla W(x)^{\top}f(x)=-a(1+W(x))(1-W(x))\|x\|^{p}$.
We refer the readers to liu2025physics for more theoretical details. This theorem enables us to obtain the ground truth value of $W$ during the training without solely depending on the PDE residual loss, making the training process more robust, and can provide better learning dynamics for the Lyapunov function.
@@ 3.1 Learning of Neural Controller and Lyapunov Function
$u_{\theta}(x)$ $=c\cdot\tanh(\mathrm{NN}_{c}(x)-\mathrm{NN}_{c}(x^{\ast})+\tanh^{-1}(\frac{u^{\ast}}{c})),\quad V_{\theta}(x)=\operatorname{sigmoid}(\mathrm{NN}_{v}(x))$ (7)
This parametrization ensures that we have $0\leq V_{\theta}\leq 1$ and the crucial property $u_{\theta}(x^{\ast})=u^{\ast}$, which ensures that the origin is also an equilibrium for the closed-loop system. Moreover, this parametrization constrains the control input lie in the range [c,c][-c,c]. Sometimes, if a non-symmetric limit $[0,c]$ is more natural, we append another ReLU layer to $u_{\theta}$. This operation can also be applied to each coordinate separately to support mixed control-input limits.
@@ 3.1.1 ROA Estimation Stage
The first stage of colearning which we call ROA estimation stage, aims to set up a mostly working controller and an estimate of the Lyapunov function that can be further refined. As we have discussed, in this stage the most important aspect is the selection of training data and training domain.
Zubov Guided Sampling: Intuitively, a good training data selection scheme should mix the data points inside and outside of the running region of attraction. Selecting only “outside” points risks collapsing the sublevel set back toward the origin, while sampling exclusively “inside” offers no signal to enlarge it. In low dimensions, a well‑chosen training box plus random sampling can approximate this mix, but as dimensionality grows, the ROA occupies a vanishing fraction of the domain, and naive draws will almost always only select the points outside of the ROA. For example, in the cartpole system, the ROA we found at the end of training only occupies less than $1\%$ of the full training domain, and this will only be more pronounced in even higher-dimensional systems, and therefore naive sampling is infeasible.
In the traditional Lyapunov function framework, it is difficult to obtain this running estimation of the ROA. Zubov theorem, on the contrary, characterizes the exact region of attraction by the sublevel set $V^{\leq 1}$. Intuitively, in our training pipeline, since our parametrization already deems that $V$ cannot exactly equal $1$, we propose picking the training data by selecting a batch of points lying in $V_{\theta}^{\leq c}$ and another batch with $V_{\theta}(x)$ as close to $1$ as possible, with $c\approx 1$. To achieve this, we select data by performing projected gradient descent (PGD) madry2017towards with the following two objectives
@@
The first interior loss will push points to the sublevel set $V_{\theta}^{\leq c}$ and the outside loss will push points further away from the region of attraction as shown in Figure 1. A detailed discussion on how to pick $c$ is presented in Appendix A.4.
@@
Dynamic Training Domain Expansion: While previous works attempt to resolve the conservativeness introduced by a fixed training domain by allowing the ROA to touch non-trivially with it, we argue that the best way is instead to dynamically expand the domain during training. We thus propose to utilize the trajectory information to dynamically update the training box. We start from a relatively small training region and periodically perform a trajectory simulation starting from the points inside the current ROA estimation $V_{\theta}^{\leq c}$ with the same $c$ in the data selection. We then expand the training region to include the farthest states that the converged trajectories can reach. To also benefit simple systems where all trajectories converge to zero in a monotone fashion, we further enlarge the box uniformly for the final training domain. If there is no convergent trajectory for many iterations, we manually enlarge the box for further exploration. This approach simultaneously resolves conservativeness in ROA posed by fixed domain and also resolves the challenging forward invariance issue. A visualization of our approach can be found in Figure 1, where the original blue training domain is expanded further to incorporate the states along the convergent trajectories. The domain updating algorithm can be found in Algorithm 2 in appendix.
Loss Function: The training loss can be easily adapted from the Zubov theorem. The objective of the training is to train the Lyapunov function and controller to encourage $V_{\theta}$ to behave more like a Zubov function. Therefore, naturally we could minimize $V_{\theta}(0)$ and train on the PDE residual of (6) and the data loss (2.4). In the handling of data loss, previous works such as wang2024actor often simulates the trajectory till convergence to compute $\int_{0}^{\infty}\|\phi(t,x)\|^{p}\,dt$. This, however, would lead to very lengthy training and add difficulty in the implementation of efficient batched operations since different trajectories have different convergence speeds. We thus propose to replace this term using a Bellman-type dynamic programming approach. We note that the following rewrite is possible:
$W(x)$ $=\tanh(a\int_{0}^{\infty}\|\phi(t,x)\|^{p}\,dt)$ (10)
$=\tanh(a\int_{0}^{T}\|\phi(t,x)\|^{p}\,dt+\tanh^{-1}(W(\phi(T,x)))).$
@@
With this approach, only simulation to time $T$ is required. In practice, we found that this term can benefit training even if a very small $T$ like $T=0.01s$ is adopted. We also utilized a similar strategy as in (wang2024actor, ) to decouple the learning of the Lyapunov function and controller using the actor-critic framework and impose the boundary condition $V_{\theta}(y)\approx 1$ for $y\in\partial(2\Omega)$ to guide the early phase learning. Moreover, to ensure that the second CEGIS stage is numerically stable, we minimize $V_{\theta}(0)$ only with a hinge-type loss. Let $\epsilon>0$ be a small positive number. The final loss function is a weighted combination of the following components:
$L_{\text{zero}}(\theta)$ $=\text{ReLU}\left(V_{\theta}^{2}(0)-\epsilon\right),$ (11)
$L_{\text{pde}}(\theta)$ $=\frac{1}{k_{1}}\sum_{i=1}^{k_{1}}\left(\nabla V_{\theta}(x_{i})^{\top}f(x_{i},\text{detach}(u(x_{i})))+a(1+V_{\theta}(x_{i}))(1-V_{\theta}(x_{i}))\|x_{i}\|^{p}\right)^{2},$ (12)
$L_{\text{data}}(\theta)$ $=\frac{1}{k_{2}}\sum_{i=1}^{k_{2}}\left(V_{\theta}(x_{i})-\tanh(a\int_{0}^{T}\|\phi(t,x_{i})\|^{p}\,dt)-\tanh^{-1}(V_{\theta}(\phi(T,x_{i})))\right)^{2},$ (13)
$L_{\text{controller}}(\theta)$ $=\frac{1}{k_{1}}\sum_{i=1}^{k_{1}}\text{detach}(\nabla V_{\theta}(x_{i}))^{\top}f(x_{i},u(x_{i})),$ (14)
$L_{\text{boundary}}(\theta)$ $=\frac{1}{k_{3}}\sum_{i=1}^{k_{3}}\left(V_{\theta}(y_{i})-1\right)^{2},$ (15)
where $k_{1},k_{2},k_{3}>0$ represents the number of data points. Here, the PDE loss minimizes the PDE residual (6) by optimizing the Lyapunov function, whereas the controller loss optimizes the controller akin to a Sontag-style formula. Here $x_{i}$ are training data sampled by our novel sampling scheme, and $y_{i}$ are the boundary points from $\partial(2\Omega)$. More details about the loss function design can be found in appendix A.3.
3.1.2 CEGIS-based Refinement